# The search for Population III: Confirmation of a HeII emitter with no metal lines at z=10.6

### 20 Mar 2026 v1

@@ 7 The Population III scenario
Fig.5 shows the EW(HeII1640) vs HeII/H diagnostic diagram. This diagram was already presented in Maiolino et al. (2024b) for Hebe. In that paper, however, H was measured along the shutter using the prism, hence likely only grazing at the more diffuse emission and not probing the core of the Hebe (see discussion in Section 6). With the new data presented in Übler et al. (2026), we have a direct measurement of H in the same region and with a similar spectral resolution as for the HeII emission. The various symbols show models from Nakajima and Maiolino (2022), with various ionization parameters and metallicities. Given the complete absence of metal lines, to avoid overcrowding the plot, we focus on low metallicity models, below 10% solar. Specifically, red stars indicate PopII models, with decreasing metallicity from dark to light, down to a metallicity of (i.e. ). Note that these models use BPASS binary stellar templates, which are already more energetic than standard stellar populations (see Nakajima and Maiolino, 2022, for details). Orange pentagons are AGN models, with a broad range of possible ionizing spectra, a broad range of ionizing parameters, and a very broad range of metallicities, from 10% solar to completely pristine. The squares show the PopIII models, with sizes indicating the adopted IMF (assuming a Salpeter slope): larger symbols are for stellar masses in the range , intermediate size symbols are for stellar masses in the range , and small symbols for stellar masses in the range . The blue squares are for gas densities of . The purple squares are for densities of and . The small blue points are PopIII models from Rusta et al. (2025): dark points are for Pure III pristine systems, while light points are for systems with some PopII contamination, but still dominated by PopIII (50%). The blue points are for gas densities of and ionization parameter , the purple points are for gas densities of and ionization parameter .
@@
We note that densities of explored in the models may appear somewhat extreme; however, high redshift studies have revealed a significant population of galaxies with gas densities in this range (e.g. Martinez et al., 2025). Therefore, it is very plausible that PopIII systems can have such high densities. The reason why PopIII models with high densities are significantly higher in EW is that these are above the critical density of the 2-photons nebular continuum, which is therefore suppressed, boosting the EW(HeII). Interestingly, the same does not happen for AGN; indeed, in these cases, the 2-photon continuum is negligible, and the UV continuum is dominated by the accretion disc. Therefore, increasing the gas density to for the AGN models does not have any tangible effect.
@@
The large golden symbols are the values measured for Hebe, both total (triangle) and for the individual components (circles), as labelled. The EW(HeII) and HeII/H ratios observed for Hebe, either total or for the individual components, are well above the expectations for PopII models. AGN models also struggle to reproduce the observed constraints, primarily because they generally predict lower EW(HeII); however, we will discuss in more detail the AGN scenario in section 10. Component C1 is the one most clearly aligned with the PopIII prediction, both for the low- and high-density models. Component C2 seems inconsistent with PopIII low-density models, but still consistent with PopIII high-density models. We shall mention that it is possible that other PopIII models, with steeper SED and higher yield of HeII photons, which for example are associated to rotating PopIII stars (e.g. Yoon et al., 2012; Murphy et al., 2021; Sibony et al., 2022; Wasserman et al., 2026b), could reproduce C2 with low density gas. These models were not considered in Nakajima and Maiolino (2022) nor in Rusta et al. (2025). An alternative possibility is that in C2 the very first generation of PopIII may have provided a first quick enrichment of dust, as expected by many models (see review by Schneider and Maiolino, 2024). Taking the extinction curve by Sun et al. (2026) (one of the few extinction curves measured at high redshift), an extinction of only would be enough to move the HeII/H flux ratio to the PopIII sequence. Depending on the dust distribution geometry, in a compact configuration such an extinction can be achieved with just a few solar masses of dust (see e.g. Madau and Maiolino, 2026).
@@
Overall, both for C1 and C2 excitation by PopIII stars appears to be the scenario most consistent with the current observational constraints on Hebe. In this scenario, the two components are two nearby clusters (within 400 pc of each other), probably at slightly different evolutionary stages (possibly with C2 being already contributed by some PopII) and/or slightly different local physical properties (different densities and ionization parameters). A more extensive comparison with more detailed models is provided in in the companion paper by Rusta et al. (2026), who reach similar and even more stringent conclusions. They also constrain the stellar mass of the PopIII in the range between and a few times .
@@ 8 Lyman-
An IFU G140M observation of GN-z11 covering the location of Hebe is available (Scholtz et al., 2024a). This observation is much shorter than those with G235M or G235H. Ly is not detected at the location of Hebe, with a upper limit of . This includes the aperture correction and is integrated over a velocity range of 150 km/s; the spectrum at the HeII location is shown in the Appendix. Also, H is 5 times weaker in Hebe compared to GN-z11, suggesting that the intrinsic Ly luminosity of this regions is lower by a comparable factor.
@@ 9 The Wolf-Rayet Scenario
One alternative interpretation of the observed HeII emission is the presence of a population of Wolf-Rayet (WR) stars. Specifically, in this scenario the broader HeII line could be tracing a population of very metal-poor WR stars. Although broader than H, the HeII line is much narrower (only 110 km/s) than typically observed in WR stars. Yet, the width of the HeII lines in the WR expanding envelopes is expected to decrease as a function of metallicity, as most of the radiation pressure generating the envelope winds happens via metal line-driven winds. It is indeed observed that WR stars in more metal-poor environments tend to have narrower HeII (Vink and de Koter, 2005; Crowther and Hadfield, 2006; Crowther et al., 2023; Castro et al., 2018). However, even these low metallicity cases present prominent nitrogen (or carbon) emission lines, as (by definition of WR) the expulsion of the outer layers exposes the inner He and N burning core (e.g. Hainich et al., 2014; Berg et al., 2024, 2025).
The WR phenomenon in the extremely metal-poor regime (. The photospheric wind velocities are expected to decrease steadily as a function of metallicity (Gräfener and Hamann, 2008; Sabhahit et al., 2023; Boco et al., 2025), but even in the most metal poor systems simulated () the expected WR winds are still at least a factor of two larger than the line width observed in Hebe.
@@
Finally, the double peaked profile of the HeII emission is not typical of WR stars, whose spectra are characterized by a Gaussian-like profile (e.g. Crowther et al., 2023), and it is not clear why the HeII WR feature should have a velocity offset relative to the H.
@@ 10 The direct collapse black hole or primordial black hole scenarios
Another interesting possibility is that Hebe is actually a very small accreting black hole. If so, given the absence of metal lines (see Rusta et al. 2026 and Übler et al. 2026 for metallicity constraints), this would be the most metal-poor accreting black hole ever discovered (see Maiolino et al., 2025, for another extreme example). Being in the vicinity of GN-z11, embedded in a strong Lyman-Werner radiation field, it would be the optimal environment for the formation of a Direct Collapse Black Hole (DCBH, e.g. Loeb and Rasio, 1994; Bromm and Loeb, 2003; Regan et al., 2017). There is no evidence for a broad component of H that would indicate the presence of a Broad Line Region. However, it could potentially be that the observed narrow H is itself tracing the BLR of a very small black hole, with very low velocity dispersion. Übler et al. (2026) infer an upper limit on the black hole mass of a few times . This would be borderline with expectations for DCBHs, but still consistent with the expectations (Ferrara et al., 2014).
Another formation channel can be that of a primordial black hole (PBH, Hawking, 1971; Carr and Hawking, 1974; Escrivà et al., 2024). These are expected to have a broad mass function, but recent constraints from the CMB suggest an upper limit of about (Matteri et al., 2025), which would more comfortably comply with the observational upper limit provided by Übler et al. (2026); although one should take into account their strong clustering expected for PBHs, which can result in a rapid growth via merging in the early Universe. In the PBH scenario, the vicinity of GN-z11 would allow the initially naked black holes to start accreting gas from the circumgalactic medium of the galaxy, as it plunges into its halo.
@@
Regardless of the potential nature and origin, an accreting black hole in such a metal-poor environment would be an extremely interesting finding. However, there are a few aspects that make these scenarios less plausible. First, the lower limits on the EW(HeII) are uncomfortably high for the AGN scenario, as illustrated in Fig.5; even stronger tension is found by Übler et al. (2026) when exploring the EW(H). Secondly, the observed HeII line profile is quite different from the H one. While BLR stratification can broaden HeII with respect to Balmer lines (e.g. Brazzini et al., 2026), HeII in AGNs is generally not strongly blueshifted, nor it present a double-peaked profile, unless the Balmer lines also do.
@@ 11 Excluding photoionization from the AGN in GN-z11
Yet another possibility is that the emission of GN-z11 is highly anisotropic. In this scenario, Hebe could be seeing a radiation much larger than our line of sight. Specifically, this could be happening if GN-z11 accreting at highly super-Eddington rate, in which case the ionizing radiation would be highly beamed in the direction perpendicular to the accretion disc (e.g. Madau and Haardt, 2024), and Hebe sees that disc nearly face on, while we are observing it close to edge on. We have explored this scenario more quantitatively by leveraging the framework presented in Madau and Maiolino (2026) for Little Red Dots and Little Blue Dots. We obtain that even in this configuration, the model falls short in explaining the HeII luminosity of Hebe by at least an order of magnitude (especially given the very conservative assumptions on the distance and property of the putative clump). Additionally, the highly anisotropic and fortuitous orientation scenario, in such a small field, would imply that there should be many other similar AGN in the sky at , pointing towards us with magnitude .
@@ 12 Summary and Conclusions
We have presented deep NIRSpec-IFU high-resolution (G235H) observations of GN-z11, a luminous galaxy at z=10.6. The detection of a HeII1640 emitting region in the halo of GN-z11 was previously reported based on medium-resolution (G235M) IFU data. The main goal of the new observations was to confirm and characterize the HeII emission in the halo of GN-z11. The main results can be summarized as follows:
HeII emission is confirmed by the G235H data, in a compact region, at the location of the previous G235M detection, 3 kpc from GN-z11. We have named this HeII emitter Hebe.
The HeII identification is independently confirmed by the detection of H in the same region by Übler at al. (2026).
@@
The HeII emission is spectrally resolved in two components; the H emission is primarily associated with one of them.
@@
The only models consistent with the observed EW(HeII) and HeII/H ratios are those involving PopIII stars. Further constraints supporting this scenario come from modelling the non-detection of metal lines, as discussed in a companion paper (Rusta et al. 2026).
@@ Appendix A Single Gaussian simultaneous HeII and H fit
Fig.8 shows the result of simultaneously fitting HeII and H with a single Gaussian. The corresponding fit on H is shown in the bottom panel. In this case , in contrast with the case of fitting with two Gaussian components (), resulting in a – a clear indication that the fit with a single Gaussian fit is much worse than the fit with two Gaussian components.
@@ Appendix B Metal lines non detection
Fig.9 show the spectrum of Hebe at the location zoomed at the wavelengths where typically the strongest metal lines are observed in normal galaxies, specifically, from top to bottom: CIII]1906,1908, OIII]1661,1666, NIII]1749 multiplet, CIV1548,1551, and NIV]1483,1486. Note that an unmasked faint open shutter/artefact affects a region near the NIV doublet (gray shaded region).

### 24 Mar 2026 v2

@@ 7 The Population III scenario
Fig.5 shows the EW(HeII1640) vs HeII/H diagnostic diagram. This diagram was already presented in Maiolino et al. (2024b) for Hebe. In that paper, however, H was measured along the shutter using the prism, hence likely only grazing at the more diffuse emission and not probing the core of the Hebe (see discussion in Section 6). With the new data presented in Übler et al. (2026), we have a direct measurement of H in the same region and with a similar spectral resolution as for the HeII emission. The various symbols show models from Nakajima and Maiolino (2022), with various ionization parameters and metallicities. Given the complete absence of metal lines, to avoid overcrowding the plot, we focus on low metallicity models, below 10% solar. Specifically, red stars indicate PopII models, with decreasing metallicity from dark to light, down to a metallicity of (i.e. ). Note that these models use BPASS binary stellar templates, which are already more energetic than standard stellar populations (see Nakajima and Maiolino, 2022, for details). Orange pentagons are AGN models, with a broad range of possible ionizing spectra, a broad range of ionizing parameters, and a very broad range of metallicities, from 10% solar to completely pristine. The squares show the PopIII models, with sizes indicating the adopted IMF (assuming a Salpeter slope): larger symbols are for stellar masses in the range , intermediate size symbols are for stellar masses in the range , and small symbols for stellar masses in the range . The blue squares are for gas densities of . The purple squares are for gas densities of and . The small blue points are PopIII models from Rusta et al. (2026): dark points are for Pure III pristine systems, while light points are for systems with some PopII contamination, but still dominated by PopIII (50%). The blue points are for gas densities of and ionization parameter , the purple points are for gas densities of and ionization parameter .
@@
We note that densities of explored in the models may appear somewhat extreme; however, high redshift studies have revealed a significant population of galaxies with gas densities in this range (e.g. Martinez et al., 2025; Moreschini et al., 2026). Therefore, it is very plausible that PopIII systems can have such high densities. The reason why PopIII models with high densities are significantly higher in EW is that these are above the critical density of the 2-photons nebular continuum, which is therefore suppressed, boosting the EW(HeII). Interestingly, the same does not happen for AGN; indeed, in these cases, the 2-photon continuum is negligible, and the UV continuum is dominated by the accretion disc. Therefore, increasing the gas density to for the AGN models does not have any tangible effect.
@@
The large golden symbols are the values measured for Hebe, both total (triangle) and for the individual components (circles), as labelled. The EW(HeII) and HeII/H ratios observed for Hebe, either total or for the individual components, are well above the expectations for PopII models. AGN models also struggle to reproduce the observed constraints, primarily because they generally predict lower EW(HeII); however, we will discuss in more detail the AGN scenario in section 10. Component C1 is the one most clearly aligned with the PopIII prediction, both for the low- and high-density models. Component C2 seems inconsistent with PopIII low-density models, but still consistent with PopIII high-density models. We shall mention that it is possible that other PopIII models, with steeper SED and higher yield of HeII photons, which for example are associated to rotating or binary PopIII stars (e.g. Yoon et al., 2012; Murphy et al., 2021; Sibony et al., 2022; Wasserman et al., 2026b; Lecroq et al., 2025), could reproduce C2 with low density gas. These models were not considered in Nakajima and Maiolino (2022) nor in Rusta et al. (2026). An alternative possibility is that in C2 the very first generation of PopIII may have provided a first quick enrichment of dust, as expected by many models (see review by Schneider and Maiolino, 2024). Taking the extinction curve by Sun et al. (2026) (one of the few extinction curves measured at high redshift), an extinction of only would be enough to move the HeII/H flux ratio to the PopIII sequence. Depending on the dust distribution geometry, in a compact configuration such an extinction can be achieved with just a few solar masses of dust (see e.g. Madau and Maiolino, 2026).
@@
Overall, both for C1 and C2 excitation by PopIII stars appears to be the scenario most consistent with the current observational constraints on Hebe. In this scenario, the two components are two nearby clusters (within 400 pc of each other), probably at slightly different evolutionary stages (possibly with C2 being already contributed by some PopII) and/or slightly different local physical properties (different densities and ionization parameters).
A more extensive comparison with more detailed models is provided in the companion paper by Rusta et al. (2026), who reach similar and even more stringent conclusions. They also constrain the total stellar mass of the PopIII star clusters in the range between and a few times . We note that is already well in the range expected for PopIII systems (e.g. Storck et al., 2025), although possibly towards the upper end, which is not unexpected given the peculiar environment. One should also take into account that there are various uncertain factors that could result in further reducing the inferred stellar mass, specifically: i) the lensing factor of the foreground galaxy is uncertain, and could result in a mass two times lower; ii) as mentioned, other PopIII models (especially rotating and/or binary PopIII stars) can result higher luminosity for a given mass, which would alos result in a reduced mass; iii) as we have spectrally resolved Hebe in two components, it is likely that also these two components are made of even lower mass systems, whose individual spectral signatures are blended within our spectral resolution.
We finally discuss that, as already mentioned in Section 3, there is another secondary peak in the G235H HeII map, located at about 0.8′′ North of GN-z11. Based on the R2700 alone, this would certainly not be significant. However, it is interesting that it is marginally detected at the same location also in the G235M map, which may support that this is a second HeII emitter. If so, this would fit a scenario in which the halo of GN-z11 has recently been subject to pristine gas accretion from the N-NE, which is resulting in PopIII star formation at various locations as a consequence of gas compression. We note that there is possibly a third peak to the NW of GN-z11 seen both in the G235H and G235M maps; however, this is weaker in G235H and is also located in the region of G235M (West of GN-z11) where there are fewer frames and possible edge effects, hence less reliable.
@@ 8 Lyman-
An IFU G140M observation of GN-z11 covering the location of Hebe is available (Scholtz et al., 2024a). This observation is much shorter than those with G235M or G235H. Ly is not detected at the location of Hebe, with a upper limit of . This includes the aperture correction and is integrated over a velocity range of 150 km/s. Also, H is 5 times weaker in Hebe compared to GN-z11, suggesting that the intrinsic Ly luminosity of this regions is lower by a comparable factor (PopIII and AGN are expected to have a similar HeII/Ly intrinsic ratio, Nakajima and Maiolino, 2022).
@@ 9 The Wolf-Rayet Scenario
One alternative interpretation of the observed HeII emission is the presence of a population of Wolf-Rayet (WR) stars. Specifically, in this scenario the broader HeII line, if not interpreted as two separate components, could be tracing a population of very metal-poor WR stars. Although broader than H, the HeII line is much narrower (only 110 km/s) than typically observed in WR stars. Yet, the width of the HeII lines in the WR expanding envelopes is expected to decrease as a function of metallicity, as most of the radiation pressure generating the envelope winds happens via metal line-driven winds. It is indeed observed that WR stars in more metal-poor environments tend to have narrower HeII (Vink and de Koter, 2005; Crowther and Hadfield, 2006; Crowther et al., 2023; Castro et al., 2018). However, even these low metallicity cases present prominent nitrogen (or carbon) emission lines, as (by definition of WR) the expulsion of the outer layers exposes the inner He and N burning core (e.g. Hainich et al., 2014; Berg et al., 2024, 2025).
The WR phenomenon in the extremely metal-poor regime (), has been studied only theoretically. The photospheric wind velocities are expected to decrease steadily as a function of metallicity (Gräfener and Hamann, 2008; Sabhahit et al., 2023; Boco et al., 2025), but even in the most metal poor systems simulated () the expected WR winds are still at least a factor of two larger than the line width observed in Hebe.
@@
Finally, the double peaked profile of the HeII emission in Hebe is not typical of WR stars, whose spectra are characterized by a Gaussian-like profile (e.g. Crowther et al., 2023), and it is not clear why the HeII WR feature should have a velocity offset relative to the H.
@@ 10 The direct collapse black hole or primordial black hole scenarios
Another interesting possibility is that Hebe is actually a very small accreting black hole. If so, given the absence of metal lines (see Rusta et al. 2026 and Übler et al. 2026 for metallicity constraints), this would be the most metal-poor accreting black hole ever discovered (see Maiolino et al., 2025, for another extreme example). Being in the vicinity of GN-z11, embedded in a strong Lyman-Werner radiation field, it would be the optimal environment for the formation of a Direct Collapse Black Hole (DCBH, e.g. Loeb and Rasio, 1994; Bromm and Loeb, 2003; Regan et al., 2017). There is no evidence for a broad component of H that would indicate the presence of a Broad Line Region. However, it could potentially be that the observed narrow H is itself tracing the BLR of a very small black hole, with very low velocity dispersion. Übler et al. (2026) infer an upper limit on the black hole mass of a few times . This would be borderline with expectations for DCBHs, but still consistent with the predictions (Ferrara et al., 2014).
Another formation channel can be that of a primordial black hole (PBH, Hawking, 1971; Carr and Hawking, 1974; Escrivà et al., 2024). These are expected to have a broad mass function, but recent constraints from the CMB suggest an upper limit of about (Matteri et al., 2025), which would more comfortably comply with the observational upper limit provided by Übler et al. (2026); although one should take into account the strong clustering expected for PBHs, which can result in a rapid growth via merging in the early Universe. In the PBH scenario, the vicinity of GN-z11 would allow the initially “naked” black holes to start accreting gas from the circumgalactic medium of the galaxy, as it plunges into its halo.
@@
Regardless of the potential nature and origin, an accreting black hole in such a metal-poor environment would be an extremely interesting finding. However, there are a few aspects that make these scenarios less plausible. First, the lower limits on the EW(HeII) are uncomfortably high for the AGN scenario, as illustrated in Fig.5; even stronger tension is found by Übler et al. (2026) when exploring the EW(H). Secondly, the observed HeII line profile is quite different from the H one. While BLR stratification can broaden HeII with respect to Balmer lines (e.g. Brazzini et al., 2026), HeII in AGNs is generally not strongly blueshifted, nor it presents a double-peaked profile, unless the Balmer lines also do.
@@ 11 Excluding photoionization from the AGN in GN-z11
Yet another possibility is that the emission of GN-z11 is highly anisotropic. In this scenario, Hebe could be seeing a radiation much larger than our line of sight. Specifically, this could be happening if GN-z11 is accreting at highly super-Eddington rate, in which case the ionizing radiation would be highly beamed in the direction perpendicular to the accretion disc (e.g. Madau and Haardt, 2024), and Hebe sees that disc nearly face on, while we are observing it close to edge on. We have explored this scenario more quantitatively by leveraging the framework presented in Madau and Maiolino (2026) for Little Red Dots and Little Blue Dots. We obtain that even in this configuration, the model falls short in explaining the HeII luminosity of Hebe by at least an order of magnitude (especially given the very conservative assumptions on the distance and property of the putative clump). Additionally, the highly anisotropic and fortuitous orientation scenario, in such a small field, would imply that there should be many other similar AGN in the sky at , pointing towards us with magnitude .
@@ 12 Summary and Conclusions
We have presented deep NIRSpec-IFU high-resolution (G235H) observations of Hebe, a HeII emitting object in the halo of the z=10.6 galaxy GN-z11, previously identified via medium resolution (G235M) IFU spectroscopy. The main results of the observations can be summarized as follows:
HeII emission is confirmed by the G235H data, in a compact region, at the location of the previous G235M detection, 3 kpc from GN-z11.
The HeII identification is independently confirmed by the detection of H in the same region by Übler et al. (2026).
@@
The HeII emission is spectrally resolved in two components separated by about 120 km/s; the H emission is primarily associated with one of them.
@@
The only models consistent with the observed EW(HeII) and HeII/H ratios are those involving PopIII stars. Further constraints supporting this scenario come from modelling the non-detection of metal lines, as discussed in a companion paper (Rusta et al., 2026).
@@
In summary, Hebe represents one of the most convincing pieces of evidence for Population III stars in the early Universe, and its properties support models of their formation and early evolution. However, equally compelling alternatives, such as a Direct Collapse Black Hole or a Primordial Black Hole, cannot be completely excluded.
@@ Appendix A Single Gaussian simultaneous HeII and H fit
Fig.8 shows the result of simultaneously fitting HeII and H with a single Gaussian, i.e. the the two lines are fitted with Gaussians with the same velocity and width. In this case , in contrast with the case of fitting with two Gaussian components (), resulting in a – a clear indication that the fit with a single Gaussian fit is much worse than the fit with two Gaussian components.
@@ Appendix B Metal lines non detection
Fig.9 shows the spectrum of Hebe zoomed at the wavelengths where typically the strongest metal lines are observed in normal galaxies, specifically, from top to bottom: CIII]1906,1908, OIII]1661,1666, NIII]1749 multiplet, CIV1548,1551, and NIV]1483,1486. Note that an unmasked faint open shutter/artefact affects a region near the NIV doublet (gray shaded region).